# A Survey on Automatic Credibility Assessment of Textual Credibility Signals in the Era of Large Language Models

### v1

@@ 10. Conclusions
With the rapid development of generative AI, the threat of misusing LLMs to hybrid operations is one of the commonly cited risks of their future development (Goldstein et al., 2023; Buchanan et al., 2021). The recent research demonstrated especially a high potential of LLMs to generate disinformation content in various global or local narratives (Vykopal et al., 2024). In this context, a large-scale ability to automatically determine the credibility of online social media content becomes even more crucial as it was in the pre-LLM era.
Potential negative impact on society and democratic values was recognized by the research community and previously led to emergence of many fake news detection approaches. At this place, we would like to emphasize that credibility assessment, while being a sibling task to fake-news detection, provides several considerable advantages, such as better and deeper explainability or wider opportunities for application in research as well as in practice.
@@
In addition, being able to reliably automatically recognize credible content, is also a way how to tackle with extensive spreading and consumption of false information. Detected high-credible content can be promoted by recommender systems or search engines, or highlighted in a user interface of social media platforms, while low-credible content can be accompanied with a low-credibility warning. With a breakdown of predicted credibility score to individual credibility indicators, end users (either media professionals or a general public) are able to explore the evidence leading to the credibility assessment or manually evaluate individual detected credibility signals and make a final assessment by themselves. Such level of explainability and pro-active human involvement in the decision process is vital and unfortunately lacks in many fake news detection approaches that commonly result into a single predicted value (commonly a binary one) with challenging or even impossible explanation caused by a black-box nature of the employed techniques (e.g., deep learning approaches).
@@
Despite these considerable advantages of credibility assessment, the current state-of-the-art research works suffer from multiple drawbacks. Most crucially, credibility assessment field can be characterized as highly fragmented. On one side, there are credibility assessment approaches that automatically detect credibility signals and aggregate them to make a final prediction about the content credibility. The utilized credibility signals are, however, mostly shallow linguistic ones (such as a number of hashtags), their automatic detection relies on simple methods (like rule- or heuristic-based techniques), and also prediction utilizes mostly basic weighting schemata. There are only very few approaches that are in line with the current state of the art (deep learning, LLMs), such as (Leite et al., 2024); or using the previous research results to detect more complex credibility signals, such as (Qureshi and Malick, 2023).
@@
On the other side, there are automatic approaches detecting various categories of credibility signals, like factuality, biases, persuasion techniques, or previously fact-checked claims. The prevalence of state-of-the-art techniques (including LLMs and various fine-tuning approaches, including PEFTs) is much higher in these works. However, such approaches remain isolated from credibility assessment, many times even not mentioning that their prediction can be considered as one of more advanced and more reliable credibility signals.
@@
To contribute in closing this undesired gap between research works, lack of interconnection of research results, as well as hindering application of the outcomes in practice, we conducted this systematic survey on automatic credibility assessment and detection of credibility signals from the NLP perspective. By collecting and describing 175 research papers, we not only systematically summarized the current state of the research in (currently fragmented) research areas, but also identified challenges and potential for future research. Our thorough analyses and discussions aim to point to an interesting avenues for future research – out of them, we would like to specifically highlight the adoption of latest highly-multilingual and multimodal LLMs. Their capabilities, currently remaining largely undiscovered, represent a way how to make a significant step forward in the area of credibility assessment. Moreover, they provide an unprecedented chance for deployment of the research results and providing novel credibility assessment tools to media professionals as well as general public. Such tools can serve both purposes – to detect false non-credible content as well as identify credible one that is worth further reading and sharing in social media environment.
@@ Acknowledgements.
This work was partially supported by the European Union under the Horizon Europe projects: vera.ai (GA No. 101070093), AI-CODE (GA No. 101135437); AI4TRUST (GA No. 101070190); by the UK’s innovation agency (Innovate UK) grant 10039055; and by EU NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the project No. 09I03-03-V03-00020.References
@@
- Al-alshaqi et al. (2024) Mohammed Al-alshaqi, Danda B. Rawat, and Chunmei Liu. 2024. Ensemble Techniques for Robust Fake News Detection: Integrating Transformers, Natural Language Processing, and Machine Learning. Sensors 24, 18 (2024). https://doi.org/10.3390/s24186062
@@
- Elmimouni et al. (2022) Houda Elmimouni, Andrea Forte, and Jonathan Morgan. 2022. Why People Trust Wikipedia Articles: Credibility Assessment Strategies Used by Readers. In Proceedings of the 18th International Symposium on Open Collaboration (¡conf-loc¿, ¡city¿Madrid¡/city¿, ¡country¿Spain¡/country¿, ¡/conf-loc¿) (OpenSym ’22). Association for Computing Machinery, New York, NY, USA, Article 9, 10 pages. https://doi.org/10.1145/3555051.3555052
@@
- Hamed et al. (2023) Suhaib Kh Hamed, Mohd Juzaiddin Ab Aziz, and Mohd Ridzwan Yaakub. 2023. A Review of Fake News Detection Approaches: A Critical Analysis of Relevant Studies and Highlighting Key Challenges Associated with the Dataset, Feature Representation, and Data Fusion. Heliyon 9, 10 (Oct. 2023), e20382. https://doi.org/10.1016/j.heliyon.2023.e20382
@@
- Lin et al. (2024) Luyang Lin, Lingzhi Wang, Xiaoyan Zhao, Jing Li, and Kam-Fai Wong. 2024. IndiVec: An Exploration of Leveraging Large Language Models for Media Bias Detection with Fine-Grained Bias Indicators. In Findings of the Association for Computational Linguistics: EACL 2024, Yvette Graham and Matthew Purver (Eds.). Association for Computational Linguistics, St. Julian’s, Malta, 1038–1050. https://aclanthology.org/2024.findings-eacl.70
@@
- Macko et al. (2024) Dominik Macko, Robert Moro, Adaku Uchendu, Ivan Srba, Jason Samuel Lucas, Michiharu Yamashita, Nafis Irtiza Tripto, Dongwon Lee, Jakub Simko, and Maria Bielikova. 2024. Authorship Obfuscation in Multilingual Machine-Generated Text Detection. arXiv:2401.07867 [cs.CL] https://arxiv.org/abs/2401.07867
@@
- Nakov et al. (2021a) Preslav Nakov, Giovanni Da San Martino, Tamer Elsayed, Alberto Barrón-Cedeño, Rubén Míguez, Shaden Shaar, Firoj Alam, Fatima Haouari, Maram Hasanain, Watheq Mansour, Bayan Hamdan, Zien Sheikh Ali, Nikolay Babulkov, Alex Nikolov, Gautam Kishore Shahi, Julia Maria Struß, Thomas Mandl, Mücahid Kutlu, and Yavuz Selim Kartal. 2021a. Overview of the CLEF-2021 CheckThat! Lab on Detecting Check-Worthy Claims, Previously Fact-Checked Claims, and Fake News. In Experimental IR Meets Multilinguality, Multimodality, and Interaction - 12th International Conference of the CLEF Association, CLEF 2021, Virtual Event, September 21-24, 2021, Proceedings (Lecture Notes in Computer Science, Vol. 12880), K. Selçuk Candan, Bogdan Ionescu, Lorraine Goeuriot, Birger Larsen, Henning Müller, Alexis Joly, Maria Maistro, Florina Piroi, Guglielmo Faggioli, and Nicola Ferro (Eds.). Springer, 264–291. https://doi.org/10.1007/978-3-030-85251-1_19
- Nakov et al. (2021b) Preslav Nakov, Husrev Taha Sencar, Jisun An, and Haewoon Kwak. 2021b. A survey on predicting the factuality and the bias of news media. arXiv preprint arXiv:2103.12506 (2021).
@@
- Pecher et al. (2024a) Branislav Pecher, Jan Cegin, Robert Belanec, Jakub Simko, Ivan Srba, and Maria Bielikova. 2024a. Fighting Randomness with Randomness: Mitigating Optimisation Instability of Fine-Tuning using Delayed Ensemble and Noisy Interpolation. arXiv:2406.12471 [cs.CL] https://arxiv.org/abs/2406.12471
@@
- Pecher et al. (2024c) Branislav Pecher, Ivan Srba, and Maria Bielikova. 2024c. On Sensitivity of Learning with Limited Labelled Data to the Effects of Randomness: Impact of Interactions and Systematic Choices. arXiv:2402.12817 [cs.CL] https://arxiv.org/abs/2402.12817
@@
- Pustejovsky et al. (2003) James Pustejovsky, Patrick Hanks, Roser Sauri, Andrew See, Robert Gaizauskas, Andrea Setzer, Dragomir Radev, Beth Sundheim, David Day, Lisa Ferro, et al. 2003. The timebank corpus. In Corpus linguistics, Vol. 2003. Lancaster, UK, 40.
@@
- Suprem et al. (2022) Abhijit Suprem, Sanjyot Vaidya, and Calton Pu. 2022. Exploring Generalizability of Fine-Tuned Models for Fake News Detection. In 2022 IEEE 8th International Conference on Collaboration and Internet Computing (CIC). IEEE Computer Society, 82–88. https://doi.org/10.1109/CIC56439.2022.00022
@@
- Wang et al. (2023) Ace Wang, Liz Maylin De Jesus Sanchez, Anya Wintner, Yuanxin Zhu, and Eni Mustafaraj. 2023. Assessing Google Search’s New Features in Supporting Credibility judgments of Unknown Websites. In Proceedings of the 2023 Conference on Human Information Interaction and Retrieval (CHIIR ’23). Association for Computing Machinery, New York, NY, USA, 303–307. https://doi.org/10.1145/3576840.3578277
@@
- Yan et al. (2024) Yeqing Yan, Peng Zheng, and Yongjun Wang. 2024. Enhancing large language model capabilities for rumor detection with Knowledge-Powered Prompting. Engineering Applications of Artificial Intelligence 133 (2024), 108259. https://doi.org/10.1016/j.engappai.2024.108259

### v2

@@ 10. Conclusions
With the rapid development of generative AI, the potential misuse of LLMs for hybrid operations has become one of the most frequently cited risks (Goldstein et al., 2023; Buchanan et al., 2021). Recent research has highlighted the particularly high capacity of LLMs to generate (personalized) disinformation across both global and local narratives (Vykopal et al., 2024; Zugecova et al., 2025). In this context, the large-scale capability to automatically assess the credibility of online social media content becomes even more crucial as it was in the pre-LLM era.
Potential negative impact on society and democratic values was recognized by the research community and previously led to emergence of many false information detection approaches. Unfortunately, automatic credibility assessment and the detection of credibility signals have not yet received the same level of attention. At this point, we emphasize that credibility assessment, while closely related to false information detection, provides complementary insights. In particular, false information that appears credible can have far more detrimental effects than content that is easily recognized as non-credible.
@@
Thanks to inherent two-step nature of credibility assessment (to detect credibility signals and then to aggregate them into a credibility label/score), credibility assessment encompasses a high level of explainability and wide opportunities for application in research as well as in practice. Detected high-credible content can be promoted by recommender systems or search engines, or highlighted in a user interface of social media platforms, while low-credible content can be accompanied with a low-credibility warning. With a breakdown of predicted credibility score to individual credibility indicators, end users (either media professionals or the general public) are able to explore the evidence leading to the credibility assessment or manually evaluate individual detected credibility signals and make a final assessment by themselves. Such level of explainability and pro-active human involvement in the decision process is vital and unfortunately lacks in many current false information detection approaches that commonly result into a single predicted value (commonly a binary one) with challenging or even impossible explanation caused by a black-box nature of the employed techniques (e.g., deep learning approaches).
@@
Despite these considerable advantages of credibility assessment, the current state-of-the-art research works suffer from multiple drawbacks. Most crucially, credibility assessment field can be characterized as highly fragmented. On one side, there are credibility assessment approaches that automatically detect credibility signals and aggregate them to make a final prediction about the content credibility. The utilized credibility signals are, however, mostly shallow linguistic ones (such as a number of hashtags), their automatic detection relies mostly on outdated and simple methods (like rule- or heuristic-based techniques), and also prediction utilizes mostly basic weighting schemata. There are only very few approaches that are in line with the current state of the art (deep learning, LLMs), such as (Leite et al., 2024); or building upon the previous research results to detect more complex credibility signals, such as (Qureshi and Malick, 2023).
@@
On the other side, there are automatic approaches detecting various categories of credibility signals, like event factuality, biases, persuasion techniques, or previously fact-checked claims. The prevalence of state-of-the-art techniques (including LLMs and various fine-tuning approaches, including PEFTs) is much higher in these works. However, such approaches remain isolated from credibility assessment, many times even not mentioning that their prediction can be considered as one of more advanced and more reliable credibility signals.
@@
To address this undesired gap between research works, lack of interconnection of research results, as well as hindering application of the outcomes in practice, we conducted this systematic survey on automatic credibility assessment and detection of credibility signals from the NLP perspective. By collecting and describing 175 research papers, we not only systematically summarised the current state of the research in (currently fragmented) research areas, but also identified shared challenges and potential for future research. Our thorough analyses and discussions aim to point to an interesting avenues for future research – particularly, we would like to highlight three most promising research streams:
Adoption of advanced multilingual and multimodal LLMs. The emerging multilingual and multimodal LLMs offer substantial – yet still largely untapped – potential for advancing credibility assessment. Their capabilities can significantly improve both detection performance and language coverage. Moreover, these models open new avenues for deploying credibility assessment tools to support both media professionals and the general public. Such tools can serve both purposes – to detect non-credible content as well as to identify credible one that is worth further reading and sharing in social media environment.
Development of multilingual and multi-category benchmark datasets. The current lack of standardized, multilingual benchmark datasets hampers direct comparison between methods and slows progress in the field. We advocate for future work to focus on unifying existing – but often fragmented – efforts into a shared benchmark that would include diverse credibility signals annotated over the same content. Such a resource would enable novel research opportunities – such as multi-task learning or novel credibility assessment algorithms – and at the same time, would potentially reduce current redundant efforts in dataset creation. Especially existing highly multilingual datasets used for previously fact-checked claim retrieval could serve as a foundation for building such a benchmark.
Addressing practical deployment challenges. Finally, bridging the gap between academic research and real-world deployment requires greater attention to practical considerations – ranging from ethical and legal implications to the robustness of models against out-of-distribution and even adversarial inputs. Additionally, issues of computational efficiency and explainability must be addressed to ensure credibility assessment systems are both effective, trustful and reliable. Despite these challenges, existing tools already demonstrate the positive real-world impact such systems can have, both in supporting media professionals and enhancing the information ecosystem for everyday users.
@@ Acknowledgements.
This work was partially supported by the European Union under the Horizon Europe projects: vera.ai (GA No. 101070093), AI-CODE (GA No. 101135437), and by AI4TRUST (GA No. 101070190); by the UK’s innovation agency (Innovate UK) grant 10039055; by EU NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the project No. 09I03-03-V03-00020.References
@@
- Augenstein et al. (2024) Isabelle Augenstein, Timothy Baldwin, Meeyoung Cha, Tanmoy Chakraborty, Giovanni Luca Ciampaglia, David Corney, Renee DiResta, Emilio Ferrara, Scott Hale, Alon Halevy, et al. 2024. Factuality challenges in the era of large language models and opportunities for fact-checking. Nature Machine Intelligence 6, 8 (2024), 852–863.
@@
- Bang et al. (2024) Yejin Bang, Delong Chen, Nayeon Lee, and Pascale Fung. 2024. Measuring political bias in large language models: What is said and how it is said. arXiv preprint arXiv:2403.18932 (2024).
@@
- Breum et al. (2024) Simon Martin Breum, Daniel Vædele Egdal, Victor Gram Mortensen, Anders Giovanni Møller, and Luca Maria Aiello. 2024. The Persuasive Power of Large Language Models. Proceedings of the International AAAI Conference on Web and Social Media 18, 1 (May 2024), 152–163. https://doi.org/10.1609/icwsm.v18i1.31304
@@
- Cegin et al. (2024) Jan Cegin, Branislav Pecher, Jakub Simko, Ivan Srba, Maria Bielikova, and Peter Brusilovsky. 2024. Effects of diversity incentives on sample diversity and downstream model performance in LLM-based text augmentation. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Lun-Wei Ku, Andre Martins, and Vivek Srikumar (Eds.). Association for Computational Linguistics, Bangkok, Thailand, 13148–13171. https://doi.org/10.18653/v1/2024.acl-long.710
@@
- Chen and Shu (2023) Canyu Chen and Kai Shu. 2023. Can llm-generated misinformation be detected? arXiv preprint arXiv:2309.13788 (2023).
@@
- Elmimouni et al. (2022) Houda Elmimouni, Andrea Forte, and Jonathan Morgan. 2022. Why People Trust Wikipedia Articles: Credibility Assessment Strategies Used by Readers. In Proceedings of the 18th International Symposium on Open Collaboration (OpenSym ’22). Association for Computing Machinery, New York, NY, USA, Article 9, 10 pages. https://doi.org/10.1145/3555051.3555052
@@
- Hellmueller and Trilling (2012) Lea Hellmueller and Damian Trilling. 2012. The Credibility of Credibility Measures: A Meta-Analysis in Leading Communication Journals, 1951 to 2011. In World Association for Public Opinion Research (WAPOR). Hongkong.
@@
- Hovland and Weiss (1951) Carl I. Hovland and Walter Weiss. 1951. The Influence of Source Credibility on Communication Effectiveness*. Public Opinion Quarterly 15, 4 (01 1951), 635–650. https://doi.org/10.1086/266350 arXiv:https://academic.oup.com/poq/article-pdf/15/4/635/5422100/15-4-635.pdf
@@
- Iqbal et al. (2024) Hasan Iqbal, Yuxia Wang, Minghan Wang, Georgi Georgiev, Jiahui Geng, Iryna Gurevych, and Preslav Nakov. 2024. OpenFactCheck: A Unified Framework for Factuality Evaluation of LLMs. arXiv preprint arXiv:2408.11832 (2024).
@@
- Jin et al. (2024) Yiping Jin, Leo Wanner, and Alexander Shvets. 2024. GPT-HateCheck: Can LLMs Write Better Functional Tests for Hate Speech Detection?. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), Nicoletta Calzolari, Min-Yen Kan, Veronique Hoste, Alessandro Lenci, Sakriani Sakti, and Nianwen Xue (Eds.). ELRA and ICCL, Torino, Italia, 7867–7885. https://aclanthology.org/2024.lrec-main.694/
@@
- Lin et al. (2024a) Luyang Lin, Lingzhi Wang, Jinsong Guo, and Kam-Fai Wong. 2024a. Investigating bias in llm-based bias detection: Disparities between llms and human perception. arXiv preprint arXiv:2403.14896 (2024).
- Lin et al. (2024b) Luyang Lin, Lingzhi Wang, Xiaoyan Zhao, Jing Li, and Kam-Fai Wong. 2024b. IndiVec: An Exploration of Leveraging Large Language Models for Media Bias Detection with Fine-Grained Bias Indicators. In Findings of the Association for Computational Linguistics: EACL 2024, Yvette Graham and Matthew Purver (Eds.). Association for Computational Linguistics, St. Julian’s, Malta, 1038–1050. https://aclanthology.org/2024.findings-eacl.70
@@
- Lunardi et al. (2024) Riccardo Lunardi, David La Barbera, and Kevin Roitero. 2024. The Elusiveness of Detecting Political Bias in Language Models. In Proceedings of the 33rd ACM International Conference on Information and Knowledge Management. 3922–3926.
@@
- Macko et al. (2025) Dominik Macko, Jakub Kopál, Robert Moro, and Ivan Srba. 2025. MultiSocial: Multilingual Benchmark of Machine-Generated Text Detection of Social-Media Texts. In Proceedings of the 63rd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Wanxiang Che, Joyce Nabende, Ekaterina Shutova, and Mohammad Taher Pilehvar (Eds.). Association for Computational Linguistics, Vienna, Austria, 727–752. https://doi.org/10.18653/v1/2025.acl-long.36
@@
- Macko et al. (2024) Dominik Macko, Robert Moro, Adaku Uchendu, Ivan Srba, Jason S Lucas, Michiharu Yamashita, Nafis Irtiza Tripto, Dongwon Lee, Jakub Simko, and Maria Bielikova. 2024. Authorship Obfuscation in Multilingual Machine-Generated Text Detection. In Findings of the Association for Computational Linguistics: EMNLP 2024, Yaser Al-Onaizan, Mohit Bansal, and Yun-Nung Chen (Eds.). Association for Computational Linguistics, Miami, Florida, USA, 6348–6368. https://doi.org/10.18653/v1/2024.findings-emnlp.369
@@
- Matz et al. (2024) Sandra C Matz, Jacob D Teeny, Sumer S Vaid, Heinrich Peters, Gabriella M Harari, and Moran Cerf. 2024. The potential of generative AI for personalized persuasion at scale. Scientific Reports 14, 1 (2024), 4692.
@@
- Mohanty (2025) Suvendu Mohanty. 2025. Fine-Grained Bias Detection in LLM: Enhancing detection mechanisms for nuanced biases. arXiv preprint arXiv:2503.06054 (2025).
@@
- Mujahid et al. (2025) Zain Muhammad Mujahid, Dilshod Azizov, Maha Tufail Agro, and Preslav Nakov. 2025. Profiling News Media for Factuality and Bias Using LLMs and the Fact-Checking Methodology of Human Experts. arXiv preprint arXiv:2506.12552 (2025).
@@
- Nakov et al. (2024) Preslav Nakov, Jisun An, Haewoon Kwak, Muhammad Arslan Manzoor, Zain Muhammad Mujahid, and Husrev Taha Sencar. 2024. A Survey on Predicting the Factuality and the Bias of News Media. (Aug. 2024), 15947–15962. https://doi.org/10.18653/v1/2024.findings-acl.944
@@
- Nakov et al. (2021) Preslav Nakov, Giovanni Da San Martino, Tamer Elsayed, Alberto Barrón-Cedeño, Rubén Míguez, Shaden Shaar, Firoj Alam, Fatima Haouari, Maram Hasanain, Watheq Mansour, Bayan Hamdan, Zien Sheikh Ali, Nikolay Babulkov, Alex Nikolov, Gautam Kishore Shahi, Julia Maria Struß, Thomas Mandl, Mücahid Kutlu, and Yavuz Selim Kartal. 2021. Overview of the CLEF-2021 CheckThat! Lab on Detecting Check-Worthy Claims, Previously Fact-Checked Claims, and Fake News. In Experimental IR Meets Multilinguality, Multimodality, and Interaction - 12th International Conference of the CLEF Association, CLEF 2021, Virtual Event, September 21-24, 2021, Proceedings (Lecture Notes in Computer Science, Vol. 12880), K. Selçuk Candan, Bogdan Ionescu, Lorraine Goeuriot, Birger Larsen, Henning Müller, Alexis Joly, Maria Maistro, Florina Piroi, Guglielmo Faggioli, and Nicola Ferro (Eds.). Springer, 264–291. https://doi.org/10.1007/978-3-030-85251-1_19
@@
- Pastorino et al. (2024) Valeria Pastorino, Jasivan A Sivakumar, and Nafise Sadat Moosavi. 2024. Decoding News Narratives: A Critical Analysis of Large Language Models in Framing Detection. arXiv preprint arXiv:2402.11621 (2024).
@@
- Pauli et al. (2024) Amalie Brogaard Pauli, Isabelle Augenstein, and Ira Assent. 2024. Measuring and Benchmarking Large Language Models’ Capabilities to Generate Persuasive Language. https://doi.org/10.48550/arXiv.2406.17753 arXiv:2406.17753.
@@
- Pecher et al. (2024a) Branislav Pecher, Jan Cegin, Robert Belanec, Jakub Simko, Ivan Srba, and Maria Bielikova. 2024a. Fighting Randomness with Randomness: Mitigating Optimisation Instability of Fine-Tuning using Delayed Ensemble and Noisy Interpolation. In Findings of the Association for Computational Linguistics: EMNLP 2024, Yaser Al-Onaizan, Mohit Bansal, and Yun-Nung Chen (Eds.). Association for Computational Linguistics, Miami, Florida, USA, 11005–11044. https://doi.org/10.18653/v1/2024.findings-emnlp.644
@@
- Pecher et al. (2024c) Branislav Pecher, Ivan Srba, and Maria Bielikova. 2024c. On Sensitivity of Learning with Limited Labelled Data to the Effects of Randomness: Impact of Interactions and Systematic Choices. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, Yaser Al-Onaizan, Mohit Bansal, and Yun-Nung Chen (Eds.). Association for Computational Linguistics, Miami, Florida, USA, 522–556. https://doi.org/10.18653/v1/2024.emnlp-main.32
@@
- Peng et al. (2025) Qiwei Peng, Robert Moro, Michal Gregor, Ivan Srba, Simon Ostermann, Marian Simko, Juraj Podrouzek, Matúš Mesarčík, Jaroslav Kopčan, and Anders Søgaard. 2025. SemEval-2025 Task 7: Multilingual and Crosslingual Fact-Checked Claim Retrieval. In Proceedings of the 19th International Workshop on Semantic Evaluation (SemEval-2025), Sara Rosenthal, Aiala Rosá, Debanjan Ghosh, and Marcos Zampieri (Eds.). Association for Computational Linguistics, Vienna, Austria, 2498–2511. https://aclanthology.org/2025.semeval-1.323/
@@
- Pisarevskaya and Zubiaga (2025) Dina Pisarevskaya and Arkaitz Zubiaga. 2025. Zero-shot and Few-shot Learning with Instruction-following LLMs for Claim Matching in Automated Fact-checking. In Proceedings of the 31st International Conference on Computational Linguistics, Owen Rambow, Leo Wanner, Marianna Apidianaki, Hend Al-Khalifa, Barbara Di Eugenio, and Steven Schockaert (Eds.). Association for Computational Linguistics, Abu Dhabi, UAE, 9721–9736. https://aclanthology.org/2025.coling-main.650/
@@
- Przybyła et al. (2024) Piotr Przybyła, Alexander Shvets, and Horacio Saggion. 2024. Verifying the robustness of automatic credibility assessment. Natural Language Processing (2024), 1–29. https://doi.org/10.1017/nlp.2024.54
- Przybyła and Soto (2021) Piotr Przybyła and Axel J. Soto. 2021. When classification accuracy is not enough: Explaining news credibility assessment. Information Processing & Management 58, 5 (2021), 102653. https://doi.org/10.1016/j.ipm.2021.102653
@@
- Ramponi et al. (2025) Alan Ramponi, Marco Rovera, Robert Moro, and Sara Tonelli. 2025. Multilingual vs Crosslingual Retrieval of Fact-Checked Claims: A Tale of Two Approaches. arXiv:2505.22118 [cs.CL] https://arxiv.org/abs/2505.22118
@@
- Rogiers et al. (2024) Alexander Rogiers, Sander Noels, Maarten Buyl, and Tijl De Bie. 2024. Persuasion with Large Language Models: a Survey. https://doi.org/10.48550/arXiv.2411.06837 arXiv:2411.06837 [cs].
- Rovera et al. (2025) Marco Rovera, Serena Cristoforetti, and Sara Tonelli. 2025. ModaFact: Multi-paradigm Evaluation for Joint Event Modality and Factuality Detection. In Proceedings of the 31st International Conference on Computational Linguistics, Owen Rambow, Leo Wanner, Marianna Apidianaki, Hend Al-Khalifa, Barbara Di Eugenio, and Steven Schockaert (Eds.). Association for Computational Linguistics, Abu Dhabi, UAE, 6378–6396. https://aclanthology.org/2025.coling-main.425/
@@
- Shokri et al. (2024) Mohammad Shokri, Vivek Sharma, Elena Filatova, Shweta Jain, and Sarah Levitan. 2024. Subjectivity Detection in English News using Large Language Models. In Proceedings of the 14th Workshop on Computational Approaches to Subjectivity, Sentiment, & Social Media Analysis, Orphée De Clercq, Valentin Barriere, Jeremy Barnes, Roman Klinger, João Sedoc, and Shabnam Tafreshi (Eds.). Association for Computational Linguistics, Bangkok, Thailand, 215–226. https://doi.org/10.18653/v1/2024.wassa-1.17
@@
- Simchon et al. (2024) Almog Simchon, Matthew Edwards, and Stephan Lewandowsky. 2024. The persuasive effects of political microtargeting in the age of generative artificial intelligence. PNAS Nexus 3, 2 (01 2024), pgae035. https://doi.org/10.1093/pnasnexus/pgae035 arXiv:https://academic.oup.com/pnasnexus/article-pdf/3/2/pgae035/56683743/pgae035.pdf
@@
- Suwaileh et al. (2025) Reem Suwaileh, Maram Hasanain, Fatema Hubail, Wajdi Zaghouani, and Firoj Alam. 2025. ThatiAR: subjectivity detection in Arabic news sentences. In Proceedings of the International AAAI Conference on Web and Social Media, Vol. 19. 2587–2602.
@@
- Vykopal et al. (2025a) Ivan Vykopal, Martin Hyben, Robert Moro, Michal Gregor, and Jakub Simko. 2025a. A Generative-AI-Driven Claim Retrieval System Capable of Detecting and Retrieving Claims from Social Media Platforms in Multiple Languages. arXiv:2504.20668 [cs.CL] https://arxiv.org/abs/2504.20668
- Vykopal et al. (2025b) Ivan Vykopal, Matúš Pikuliak, Simon Ostermann, Tatiana Anikina, Michal Gregor, and Marián Šimko. 2025b. Large Language Models for Multilingual Previously Fact-Checked Claim Detection. arXiv:2503.02737 [cs.CL] https://arxiv.org/abs/2503.02737
@@
- Wang et al. (2023a) Ace Wang, Liz Maylin De Jesus Sanchez, Anya Wintner, Yuanxin Zhu, and Eni Mustafaraj. 2023a. Assessing Google Search’s New Features in Supporting Credibility judgments of Unknown Websites. In Proceedings of the 2023 Conference on Human Information Interaction and Retrieval (CHIIR ’23). Association for Computing Machinery, New York, NY, USA, 303–307. https://doi.org/10.1145/3576840.3578277
- Wang et al. (2023b) Cunxiang Wang, Xiaoze Liu, Yuanhao Yue, Xiangru Tang, Tianhang Zhang, Cheng Jiayang, Yunzhi Yao, Wenyang Gao, Xuming Hu, Zehan Qi, et al. 2023b. Survey on factuality in large language models: Knowledge, retrieval and domain-specificity. arXiv preprint arXiv:2310.07521 (2023).
@@
- Wang et al. (2024c) Yuxia Wang, Minghan Wang, Hasan Iqbal, Georgi Georgiev, Jiahui Geng, and Preslav Nakov. 2024c. Openfactcheck: A unified framework for factuality evaluation of llms. arXiv preprint arXiv:2405.05583 (2024).
@@
- Wu et al. (2025) Junchao Wu, Shu Yang, Runzhe Zhan, Yulin Yuan, Lidia Sam Chao, and Derek Fai Wong. 2025. A Survey on LLM-Generated Text Detection: Necessity, Methods, and Future Directions. Computational Linguistics 51, 1 (March 2025), 275–338. https://doi.org/10.1162/coli_a_00549
@@
- Xu et al. (2024) Rongwu Xu, Brian S. Lin, Shujian Yang, Tianqi Zhang, Weiyan Shi, Tianwei Zhang, Zhixuan Fang, Wei Xu, and Han Qiu. 2024. The Earth is Flat because…: Investigating LLMs’ Belief towards Misinformation via Persuasive Conversation. https://doi.org/10.48550/arXiv.2312.09085 arXiv:2312.09085.
@@
- Yang et al. (2020) Fumeng Yang, Zhuanyi Huang, Jean Scholtz, and Dustin L. Arendt. 2020. How do visual explanations foster end users’ appropriate trust in machine learning?. In Proceedings of the 25th International Conference on Intelligent User Interfaces (Cagliari, Italy) (IUI ’20). Association for Computing Machinery, New York, NY, USA, 189–201. https://doi.org/10.1145/3377325.3377480
@@
- Zhang et al. (2019) Yifan Zhang, Giovanni Da San Martino, Alberto Barrón-Cedeño, Salvatore Romeo, Jisun An, Haewoon Kwak, Todor Staykovski, Israa Jaradat, Georgi Karadzhov, Ramy Baly, Kareem Darwish, James Glass, and Preslav Nakov. 2019. Tanbih: Get To Know What You Are Reading. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP): System Demonstrations, Sebastian Padó and Ruihong Huang (Eds.). Association for Computational Linguistics, Hong Kong, China, 223–228. https://doi.org/10.18653/v1/D19-3038
@@
- Zugecova et al. (2025) Aneta Zugecova, Dominik Macko, Ivan Srba, Robert Moro, Jakub Kopál, Katarína Marcinčinová, and Matúš Mesarčík. 2025. Evaluation of LLM Vulnerabilities to Being Misused for Personalized Disinformation Generation. In Proceedings of the 63rd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Wanxiang Che, Joyce Nabende, Ekaterina Shutova, and Mohammad Taher Pilehvar (Eds.). Association for Computational Linguistics, Vienna, Austria, 780–797. https://doi.org/10.18653/v1/2025.acl-long.38